OpenStudy (anonymous):

A cyclist rode 40 miles before having a flat tire and then walking 5 miles to a service station. the cyclist rate was four times the walking rate. the time spent cycling and walking was 5 hours. find the rate at which the cyclist was riding

OpenStudy (dumbcow):

d=r*t 40=C*t1 5=W*t2 C=4*W t1+t2=5 solve for C